# Asymptotic regularity of a generalised stochastic Halpern scheme with applications

### v1

@@ Abstract.
We provide abstract, general and highly uniform rates of asymptotic regularity for a generalized stochastic Halpern-style iteration, which incorporates a second mapping in the style of a Krasnoselskii-Mann iteration. This iteration is general in two ways: First, it incorporates stochasticity in a completely abstract way rather than fixing a sampling method; secondly, it includes as special cases stochastic versions of various schemes from the optimization literature, including Halpern’s iteration as well as a Krasnoselskii-Mann iteration with Tikhonov regularization terms in the sense of Boţ, Csetnek and Meier. For these particular cases, we in particular obtain linear rates of asymptotic regularity, matching (or improving) the currently best known rates for these iterations in stochastic optimization, and quadratic rates of asymptotic regularity are obtained in the context of inner product spaces for the general iteration. We utilize these rates to give bounds on the oracle complexity of such iterations under suitable variance assumptions and batching strategies, again presented in an abstract style. Finally, we sketch how the schemes presented here can be instantiated in the context of reinforcement learning to yield novel methods for Q-learning.
a Department of Mathematics, Technische Universität Darmstadt,
Schlossgartenstraße 7, 64289 Darmstadt, Germany,
b Department of Computer Science, University of Bath,
@@
E-mails: pischke@mathematik.tu-darmstadt.de, trjp20@bath.ac.uk
Keywords: Asymptotic regularity, Halpern iteration, Tikhonov regularization, Q-learning, proof mining
@@ 1. Introduction
Approximating fixed points of nonexpansive mappings is one of the most fundamental tasks in nonlinear analysis and optimization. The problem becomes particularly interesting when we only have noisy versions those mappings, in which case the resulting approximation methods become stochastic processes. Concrete examples of this general situation include model-free reinforcement learning algorithms, where variants of $Q$-learning, for instance, can be viewed as stochastic methods for computing fixpoints of nonexpansive operators.
To be more concrete, let $(X,\left\lVert\cdot\right\rVert)$ be a seperable real-valued normed space and $T,U:X\to X$ be two nonexpansive mappings on $X$, i.e.
@@
The schema (1) is designed to capture situations in which one does not have direct access to $Tx_{n}$ (respectively $Uy_{n}$), but can only use noisy versions $\tilde{T}$ of $T$ and $\tilde{U}$ of $U$. Intuitively, $\xi_{n}$ would then represent the difference between $Tx_{n}$ and the corresponding approximation of $Tx_{n}$ obtained from $\tilde{T}$ through a suitable sampling method (and similarly for $\delta_{n}$ and $Uy_{n}$), though our presentation is fully abstract and we will make no assumptions about $\xi_{n}$ or $\delta_{n}$, other than imposing controls on the means $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta\right\rVert_{n}]$, which however are carefully chosen so that they can be achieved in concrete scenarios through sampling methods such as minibatching under suitable assumptions.
@@
With this in mind, our main scheme (1) represents a stochastic analogue of the deterministic Halpern-Mann scheme recently introduced in [13], which integrates two of the most prominent methods for approximating fixed points for nonexpansive mappings, the Krasnoselskii-Mann method [23, 31] and Halpern’s method [15] (see also [40] and [41]), with the intended gain of combining the beneficial features of both, in particular the strong convergence of Halpern’s method even in infinite dimensional spaces.
@@
While all of our rates are novel, we emphasise that the qualitative asymptotic regularity results for any iteration contained in (1) that goes beyond the Halpern iteration, are also to the best of our knowledge new to the literature.
@@
The present work departs from the aforementioned case studies in nonlinear analysis in that it incorporates, for the first time, stochasticity. In this way it forms part of a recent advance of proof mining into probability theory, which comprises both new developments in the logical foundations of probability theory due to first author and Neri [32], together with applied results on the quantitative aspects of stochastic processes by the second author and Neri [33, 34]. In particular, the present paper is one of the first applications of proof mining to stochastic optimization: It represents a particularly interesting case study in this respect, in that it does not readily follow from analogous quantitative results in the deterministic setting (such as in [13, 29]), but requires a substantial arsenal of new quantitative ideas for this stochastic setting, as will be introduced and discussed in detail throughout the paper. We stress that while this logical perspective was crucial in obtaining the present results, the paper does not rely on any notions from logic at all.
The main motivation of this paper is not so much its novelty within the proof mining program, but the real applicability of the main results in concrete areas. For this reason, our last section is dedicated to reinforcement learning, where a novel method of Q-learning with Tikhonov regularization terms is sketched, and the asymptotic regularity results from the main theoretical part of this paper are utilized to give very good oracle complexity estimates for a specific mini-batching procedure that matches those obtained for Halpern’s iteration in [4]. We also emphasise that potential applications of the stochastic Halpern-Mann method are not limited to this, and in particular the presence of the second mapping in the schema (1) allows for novel instantiations in style of double Q-learning [16] that will be explored, together with the Q-learning method with Tikhonov regularization terms mentioned before, in a subsequent paper in more detail. Another promising application is represented by stochastic versions of the splitting methods with Tikhonov regularization terms as studied in [1], but we leave this for future work, as we do the study of the convergence of the method (1), where we expect that strong convergence results hold for this method also in infinite dimensional spaces, analogously to the deterministic case.
@@ 2. Preliminaries and basic lemmas
An invertible and decreasing rate $\varphi$ immediately implies the asymptotic estimate $a_{n}\leq\varphi^{-1}(n)$ for all $n\in\mathbb{N}$ and this immediately implies a complexity bound on the sequence in terms of the commonly used big O notation, namely $(a_{n})=O(\varphi^{-1}(n))$.
@@ 3.2. Asymptotic regularity relative to the mappings
Assume that
$\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$
with rate $\varphi$, and $\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\mathbb{E}[\left\lVert\delta_{n}\right\rVert],\alpha_{n}\to 0$ with rates $\rho_{1}$$\rho_{3}$, respectively. Assume further that $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with a rate $\varphi_{0}$. Then
@@
We then can similar give rates of asymptotic regularity almost surely under a slight extension of the previous conditions on the errors:
@@
Under the assumption (Hyp), assume that
$\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$
almost surely with rate $\psi$, and $\alpha_{n}\to 0$ with rate $\rho$. Further, assume that $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely with rates $\phi_{1},\phi_{2}$, respectively, and that $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with a rate $\psi_{0}$. Then
@@
If we assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively, then we can immediately derive rates $\phi_{1},\phi_{2}$ for $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely: Since $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$ with rate $\chi_{1}$, we get
@@
with $\varphi_{K,\theta,\chi}$ defined as in Theorem 3.2 and $\rho_{1},\rho_{3}$ as in Theorem 3.4. Note that this generalises known rates in this case [4, Theorem 3.3], which apply only to specific choices of the parameters. In a similar way, we get a new rate for $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely, though we do not spell it out here.
@@
for all $\lambda>0$. Let now $\lambda,\varepsilon>0$ be given. Take then $\omega$ such that $\left\lVert y_{n}(\omega)-y_{n+1}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\psi(\lambda/3,\Lambda\varepsilon/4)$ and $\left\lVert\delta_{n}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\phi(\lambda/3,\Lambda\varepsilon/4)$ as well as $\left\lVert Uy_{n+1}(\omega)-u(\omega)\right\rVert\leq\frac{K_{0}}{\lambda}$ for all $n$. Then for $n\geq\zeta(\lambda,\varepsilon)-1$, it follows from the above inequality that
$\left\lVert Uy_{n+1}(\omega)-y_{n+1}(\omega)\right\rVert<\varepsilon.$
Therefore we immediately have
$\mathbb{P}\left(\exists n\geq\zeta(\lambda,\varepsilon)\left(\left\lVert Uy_{n}-y_{n}\right\rVert\geq\varepsilon\right)\right)<\lambda$
which completes the proof. ∎
@@ 3.2.2. The general case
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Let $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with rate $\Delta$. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Lastly, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
@@
$\qquad\geq\lambda+(1-\lambda/3)+(1-\lambda/3)-2$
$\qquad=\lambda/3.$
@@
Writing $n_{0}$ for $n(\omega)$, we then have $\left\lVert Uy_{n_{0}}(\omega)-y_{n_{0}}(\omega)\right\rVert\leq 2\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq 2K^{\prime}$ so that
$\varepsilon/2\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}.$
Also, we have $\left\lVert Uy_{n_{0}}(\omega)-p\right\rVert\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}$ as well as
@@
Using a slightly different argument first devised in [18, Theorem 3.4] (see also [25, Remark 15] or [29, Remark 3.7] for similar remarks in the context of nonlinear spaces), we can slightly optimize the above rate in the context of moduli of uniform convexity of a special form: Let $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing. Then above rate $\Gamma$ holds even with $\widehat{\varepsilon}$ defined as
$\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime}).$
@@
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Under the assumptions of Theorem 3.2, let $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with rate $\Delta$ from Theorem 3.2. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Lastly, let $\mu$ be a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. Then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
@@
Before moving on to the case of fast rates for special choices of scalars and errors, we briefly discuss some natural conditions under which the previous additional assumption of a modulus of uniform integrability for the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ can be obtained (cf. [14, Chapter 5.4]), and indicate how a modulus arises from these.
@@
Another growth condition commonly imposed is to assume uniform finite expectation of the random variables under a supercoercive function. Then, in terms of an upper bound $K$ witnessing that property, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[g(\left\lVert Uy_{n}-y_{n}\right\rVert)]<K$ where $g:[0,\infty)\to[0,\infty)$ is supercoercive, i.e.
$\frac{g(x)}{x}\to\infty\text{ as }x\to\infty,$
with a rate of divergence $\kappa:(0,\infty)\to(0,\infty)$, i.e.
@@
for $a:=2K/\varepsilon$ and thus a modulus of uniform integrability is now given by
$\mu(\varepsilon):=\frac{\varepsilon}{2\kappa\left(2K/\varepsilon\right)}.$

### v2

@@ Abstract.
We provide abstract, general and highly uniform rates of asymptotic regularity for a generalized stochastic Halpern-style iteration, which incorporates a second mapping in the style of a Krasnoselskii-Mann iteration. This iteration is general in two ways: First, it incorporates stochasticity in a completely abstract way rather than fixing a sampling method; secondly, it includes as special cases stochastic versions of various schemes from the optimization literature, including Halpern’s iteration as well as a Krasnoselskii-Mann iteration with Tikhonov regularization terms in the sense of Boţ, Csetnek and Meier. For these specific cases, we in particular obtain linear rates of asymptotic regularity, matching (or improving) the currently best known rates for these iterations in stochastic optimization, and quadratic rates of asymptotic regularity are obtained in the context of inner product spaces for the general iteration. At the end, we briefly sketch how the schemes presented here can be instantiated in the context of reinforcement learning to yield novel methods for Q-learning.
Funding: The first author was supported by the ‘Deutsche Forschungsgemeinschaft’ Project DFG KO 1737/6-2. The second author was supported by the EPSRC grant EP/W035847/1.
Conflicts of interests: The authors have no relevant financial or non-financial interests to disclose.
Data availability: Data sharing not applicable to this article as no data sets were generated or analyzed.
Department of Computer Science, University of Bath,
@@
[…]
Keywords: Asymptotic regularity, Halpern iteration, Tikhonov regularization, proof mining
@@ 1. Introduction
Approximating fixed points of nonexpansive mappings is one of the most fundamental tasks in nonlinear analysis and optimization. The problem becomes particularly interesting when we only have noisy versions of those mappings, in which case the resulting approximation methods become stochastic processes. Concrete examples of this general situation include model-free reinforcement learning algorithms, where variants of $Q$-learning, for instance, can be viewed as stochastic methods for computing fixed points of nonexpansive operators.
To be more concrete, let $(X,\left\lVert\cdot\right\rVert)$ be a separable real-valued normed space and $T,U:X\to X$ be two nonexpansive mappings on $X$, i.e.
@@
Our main scheme (1) represents a stochastic analogue of the deterministic Halpern-Mann scheme for two operators, schematically appearing already in [37], and recently rediscovered in a much more generalized setting in [14]. This scheme integrates two of the most prominent methods for approximating fixed points for nonexpansive mappings, the Krasnoselskii-Mann method [24, 32] and Halpern’s method [16] (see also [42] and [43]), with the intended gain of combining the beneficial features of both, in particular the strong convergence of Halpern’s method even in infinite dimensional spaces. We conjecture that this combination gains further significance in the stochastic setting, where we anticipate that it can be used in particular to devise novel reinforcement learning algorithms or stochastic splitting methods.
In contrast to the deterministic scheme, (1) is designed to capture situations in which one does not have direct access to $Tx_{n}$ (respectively $Uy_{n}$), but can only use noisy versions $\tilde{T}$ of $T$ (and $\tilde{U}$ of $U$). Intuitively, $\xi_{n}$ would then represent the difference between $Tx_{n}$ and the corresponding approximation of $Tx_{n}$ obtained from $\tilde{T}$ through a suitable sampling method (and similarly for $\delta_{n}$ and $Uy_{n}$), though our presentation is fully abstract and we will make no assumptions about $\xi_{n}$ or $\delta_{n}$, other than imposing controls on the means $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$, which however are carefully chosen so that they can be achieved in concrete scenarios through sampling methods such as minibatching under suitable assumptions. In particular, we make no independence assumptions on the noise terms $\xi_{n}$ and $\delta_{n}$.
@@
While all of our rates are novel, we emphasise that the “qualitative” asymptotic regularity results, namely convergence alone even without considering the quantitative aspects, are also to the best of our knowledge new, for any stochastic iteration contained in (1) that goes beyond the Halpern iteration.
@@
The present work departs from the aforementioned case studies in nonlinear analysis in that it incorporates, for the first time, stochasticity. In this way it forms part of a recent advance of proof mining into probability theory, which comprises both new developments in the logical foundations of probability theory due to first author and Neri [33], together with applied results on the quantitative aspects of stochastic processes by the authors and Neri [34, 35, 36]. In particular, the present paper is one of the first applications of proof mining to stochastic optimization: It represents a particularly interesting case study in this respect, in that it does not readily follow from analogous quantitative results in the deterministic setting (such as in [14, 30]), but requires a substantial arsenal of new quantitative ideas for this stochastic setting, as will be introduced and discussed in detail throughout the paper. We stress that while this logical perspective was crucial in obtaining the present results, the paper does not rely on any notions from logic at all.
However, the main motivation of this paper is not so much its novelty within the proof mining program, but the real applicability of the main results in concrete areas. For this reason, our last section is dedicated to a brief outlook on further results arising from the present work, in particular concerning oracle complexity and reinforcement learning, and in the context of the latter how present schemes can yield novel methods for Q-learning, all of which will be discussed in further detail a forthcoming paper. While not discussed there, we also want to mention that another promising application is represented by stochastic versions of the splitting methods with Tikhonov regularization terms as studied in [2], but we leave this for future work, as we do the study of the convergence of the method (1), where we expect that strong convergence results hold for this method also in infinite dimensional spaces, analogously to the deterministic case.
@@ 2. Preliminaries and basic lemmas
The immediate benefit of such a type of rate $\varphi$ is that if it is invertible and decreasing, then we can even derive the non-asymptotic estimate $a_{n}<\varphi^{-1}(n)$ for all $n\in\mathbb{N}$, which of course further implies a complexity bound on the sequence in terms of the commonly used big O notation, namely $(a_{n})=O(\varphi^{-1}(n))$.
@@ 3.2. Asymptotic regularity relative to the mappings
Assume that $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate $\varphi$, and $\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\mathbb{E}[\left\lVert\delta_{n}\right\rVert],\alpha_{n}\to 0$ with rates $\rho_{1}$$\rho_{3}$, respectively. Assume further that $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with a rate $\varphi_{0}$. Then
@@
We then can similarly give rates of asymptotic regularity almost surely under a slight extension of the previous conditions on the errors:
@@
Under the assumption (Hyp), assume that $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate $\psi$, and $\alpha_{n}\to 0$ with rate $\rho$. Further, assume that $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely with rates $\phi_{1},\phi_{2}$, respectively, and that $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with a rate $\psi_{0}$. Then
@@
If we assume that $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$, $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively, then we can immediately derive rates $\phi_{1},\phi_{2}$ for $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely: Since $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$ with rate $\chi_{1}$, we get
@@
with $\varphi_{K,\theta,\chi}$ defined as in Theorem 3.2 and $\rho_{1},\rho_{3}$ as in Theorem 3.4. Note that this generalises known rates in this case [5, Theorem 3.3], which apply only to specific choices of the parameters. In a similar way, we get a new rate for $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely, though we do not spell it out here.
@@
for all $\lambda>0$. Let now $\lambda,\varepsilon>0$ be given. Take then $\omega$ such that $\left\lVert y_{n}(\omega)-y_{n+1}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\psi(\lambda/3,\Lambda\varepsilon/4)$ and $\left\lVert\delta_{n}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\phi(\lambda/3,\Lambda\varepsilon/4)$ as well as $\left\lVert Uy_{n+1}(\omega)-u(\omega)\right\rVert\leq\frac{K_{0}}{\lambda}$ for all $n$. Then for $n\geq\zeta(\lambda,\varepsilon)-1$, it follows from the above inequality that $\left\lVert Uy_{n+1}(\omega)-y_{n+1}(\omega)\right\rVert<\varepsilon$. This immediately yields $\mathbb{P}\left(\exists n\geq\zeta(\lambda,\varepsilon)\left(\left\lVert Uy_{n}-y_{n}\right\rVert\geq\varepsilon\right)\right)<\lambda$ which completes the proof. ∎
@@ 3.2.2. The general case
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Let $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with rate $\Delta$. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$, $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Lastly, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
@@
$\qquad\geq\lambda+(1-\lambda/3)+(1-\lambda/3)-2=\lambda/3.$
@@
Writing $n_{0}$ for $n(\omega)$, we then have $\left\lVert Uy_{n_{0}}(\omega)-y_{n_{0}}(\omega)\right\rVert\leq 2\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq 2K^{\prime}$ so that $\varepsilon/2\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}$. Also, we have $\left\lVert Uy_{n_{0}}(\omega)-p\right\rVert\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}$ as well as
@@
Using a slightly different argument first devised in [19, Theorem 3.4] (see also [26, Remark 15] or [30, Remark 3.7] for similar remarks in the context of nonlinear spaces), we can slightly optimize the above rate in the context of moduli of uniform convexity of a special form: Let $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing. Then above rate $\Gamma$ holds even with $\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime})$.
@@
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Under the assumptions of Theorem 3.2, let $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with rate $\Delta$ from Theorem 3.2. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$, $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Lastly, let $\mu$ be a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. Then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
@@
Before moving on to the case of fast rates for special choices of scalars and errors, we briefly discuss some natural conditions under which the previous additional assumption of a modulus of uniform integrability for the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ can be obtained (cf. [15, Chapter 5.4]), and indicate how a modulus arises from these.
@@
Another growth condition commonly imposed is to assume uniform finite expectation of the random variables under a supercoercive function. Then, in terms of an upper bound $K$ witnessing that property, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[g(\left\lVert Uy_{n}-y_{n}\right\rVert)]<K$ where $g:[0,\infty)\to[0,\infty)$ is supercoercive, i.e. $g(x)/x\to\infty$ as $x\to\infty$, with a rate of divergence $\kappa:(0,\infty)\to(0,\infty)$, i.e.
@@
for $a:=2K/\varepsilon$ and thus a modulus of uniform integrability is now given by $\mu(\varepsilon):=\varepsilon/2\kappa\left(2K/\varepsilon\right)$.